OpenStudy (anonymous):

Find the rate of change of elvation when X=2 f(x)=0.04(8x-x2)

OpenStudy (anonymous):

df(x)/dx = d(.04 * 8 * x)/dx + d(.04 * x^2)/dx = (.04 * 8 * 1) + (.04 * 2x) = .32 + .08x df(2)/dx = .32 + .08 * 2 = .48